# A Siamese Network to Detect If Two Iris Images Are Monozygotic

### v2

@@ Abstract
In Daugman-style iris recognition, the textures of the left and right irises of the same person are traditionally considered as being as different as the irises of two unrelated persons. However, previous research indicates that humans can detect that two iris images are from different eyes of the same person, or eyes of monozygotic twins, with an accuracy of about 80%. In this work, we employ a Siamese network architecture and contrastive learning to categorize a pair of iris images as coming from monozygotic or non-monozygotic irises. This could potentially be applied, for example, as a fast, noninvasive test to determine if twins are monozygotic or non-monozygotic. We construct a dataset comprising both synthetic monozygotic pairs (images of different irises of the same individual) and natural monozygotic pairs (images of different images from persons who are identical twins), in addition to non-monozygotic pairs from unrelated individuals, ensuring a comprehensive evaluation of the model’s capabilities. To gain deeper insights into the learned representations, we train and analyze three variants of the model using (1) the original input images, (2) iris-only images (masking everything but the iris region), and (3) non-iris-only images (masking the iris region). This comparison reveals the critical importance of iris-specific textural details and contextual ocular cues in identifying monozygotic iris patterns. The results demonstrate that models leveraging full eye-region information outperform those trained solely on iris-only data, emphasizing the nuanced interplay between iris and ocular characteristics. Our approach achieves accuracy levels using the full iris image that exceed those previously reported for human classification of monozygotic iris pairs. This study presents the first classifier designed to determine whether a pair of iris images originates from monozygotic individuals.  [^1] Code is available on https://github.com/yongleyuan/mz-iris-pairing
@@ I INTRODUCTION
Iris recognition has emerged as a leading biometric identification technology due to the unique and stable patterns present in the human iris [6]. These intricate patterns are formed randomly during fetal development [27]. Iris recognition technologies are able to identify the distinctiveness between irises from different individuals. Their high accuracy, combined with the speed and non-intrusive nature, makes them ideal for applications in security, border control, and identity verification. For example, India has used iris as one of the biometrics in the unique identification system called “Aadhaar” [24].
While the principle that “no two irises are the same” supports the effectiveness of iris recognition technology [6], a notable limitation arises when dealing with monozygotic twins. Humans can often perceive subtle similarities in the iris patterns of monozygotic (“identical”) twins, yet iris biometric systems generally fail to detect these nuances [14]. Monozygotic twins share identical genetic material due to originating from a single fertilized egg, leading to similarities in many visual characteristics, including iris patterns [17].
The ability to accurately determine if two individuals are monozygotic twins using iris recognition is crucial for several reasons. First, it offers a cost-effective and non-invasive alternative to genetic testing for distinguishing identical twins. Second, investigating how iris recognition systems detect — or fail to detect — subtle texture similarities contributes to explainable artificial intelligence (AI), improving our understanding of the decision-making processes of these systems and increasing their trustworthiness.
Problem Definition
The objective of this research is to investigate and enhance the capability of neural network-based iris recognition systems to distinguish between monozygotic and non-monozygotic twins. This involves:
@@ III-D Dataset Construction
We would like to use MZ twins for both training and testing. However, due to an insufficient number of available MZ twin samples, we use synthetic MZ iris pairs for training and actual MZ twin iris data for testing.
@@
The training and validation data used in our work will be made available to the research community on the website of the Computer Vision Research Lab at Notre Dame [^2] https://cvrl.nd.edu/projects/data/. We do not have the rights to re-distribute the MZ twins iris images [18] now, but the list of image pairs used will be made available.
@@ IV METHODOLOGY
To implement the contrastive learning strategy, we utilize a Siamese network architecture [3]. A Siamese network consists of two identical sub-networks with shared weights and parameters, ensuring both inputs are processed in the same manner. Each sub-network functions as an encoder to map an input image to a fixed-dimensional embedding in the vector space. The similarity between the two embeddings is then evaluated using a distance metric, such as Euclidean distance. The Siamese network is particularly well-suited for similarity-based tasks, as it enforces consistency in the representation of both inputs and enables the network to focus on learning meaningful differences between MZ and NMZ pairs.
@@ V-B Model Backbone
We employ ResNet-18 as the backbone for our image encoder due to its proven effectiveness in various computer vision tasks [13]. ResNet-18 is a convolutional neural network (CNN) known for its residual connection, which addresses the vanishing gradient problem by allowing gradients to flow through shortcut connections that skip one or more layers. ResNet has multiple common variants, of which ResNet-18 is the shallowest. It is effective and easy to train. We explored deeper ResNet backbones (34, 50, 101) but found that they did not offer any clear increase in accuracy for this problem; see the Supplementary Material.
@@ V-C2 Loss Function
where $\mathcal{L}$ is the loss, $y_{\text{pred}}$ is a binary label indicating whether the pair is similar/positive/MZ ($y_{\text{pred}}=1$) or dissimilar/negative/NMZ ($y_{\text{pred}}=0$), $D$ is the Euclidean distance in (1), and $m$ is the margin that defines the maximum acceptable distance between negative pairs. We set $m=1$ to normalize the distance between 0 and 1 after training so that we can easily recognize outliers if there are any. Since the distance should be normalized, we intuitively set the distance decision boundary/threshold to the midpoint: 0.5. We will discuss how this threshold affects the training and testing in Appendix -B1.
@@ V-C4 Evaluation Metrics
  • Receiver Opereating Characteristic (ROC) curve: Plots true positive rate against false positive rage.
  • Area under ROC curve (AUROC): A normalized metric that provides an aggregate measure of performance across all classifications.
@@ V-D Model Training
The model is trained separately using three different types of inputs: (1) original iris images, (2) iris-only images, and (3) non-iris-only images. The non-iris-only images function as a control group to provide insights on what areas the model is focusing on.
30% of the training data are randomly selected as validation data. These image pairs are not inputted in any forward passes. Instead, they are used after each forward pass to compute all the evaluation metrics, which will be compared with those collected after inference.
We initialized the ResNet-18 backbone with the ImageNet pre-trained weights, which should already have some prior knowledge in distinguishing between different objects. Each time the model is trained with a batch size of 32 pairs. The loss function and the optimizer, as mentioned in Sec. V-C2 and V-C3, are kept the same throughout. All experiments are run 5 times, with each individual run having the same train-validation split and data batching.
@@ V-E Training Results
Across all three types of input images, the model demonstrated rapid convergence during training. Fig. 5 illustrates accuracy and loss during training. The training accuracy for all three different input types quickly approaches 1, indicating that the model can distinguish between MZ and NMZ pairs in nearly all training samples after epoch 7. The minimal separation of lines suggests that the input type has a negligible impact on final training accuracy. The training loss follows a similar pattern of convergence, reaching near-zero around epoch 7. The iris-only image pairs show a faster decline in loss, suggesting that the iris information makes the model converge faster and easier under the same setting, compared to analyzing complete iris images and images without iris texture.
@@ V-F Testing Results
Table I presents the test outcomes for all three input types over five experimental runs. The model using original iris images achieves the best overall performance, exceeding the other two in terms of accuracy, recall, and F1 score. Although the non-iris-only input obtains the highest precision score, the advantage is marginal. Surprisingly, the model trained exclusively on iris data performs the worst. Fig. 6 shows examples of false positive and false negative predictions for all three models.
@@ VI-A Results Analysis
The testing results indicate that the original input model, which utilizes the complete eye region, outperforms both the iris-only (masked) and the non-iris-only (inversely masked) models. At first, we anticipated the masked input model to excel by focusing exclusively on the iris’s unique biometric features. However, these results suggest that non-iris ocular features or contextual information present in the original images may contribute positively to distinguishing MZ from NMZ pairs, at least under the given training conditions.
@@
  • Contextual Cues: The additional features surrounding the iris—such as sclera patterns, eyelid contours, or subtle lighting variations—could inadvertently help the model form more robust embeddings. Even though these cues are not traditionally considered primary biometric traits, they may offer supplemental distinguishing information.
@@
As for the model with non-iris-only input, while it does not outperform the model with original input, its reasonable performance suggests that even non-iris regions can sometimes aid in classification. However, it still lags behind the original model, implying that truly critical discriminative signals remain primarily in or near the iris region.
@@ VI-B Pupil Dilation Effects
VI-C Comparison with human viewer performance
@@ VII FUTURE WORKS
Systematic ablation studies examining which layers and filters contribute most to MZ/NMZ discrimination would clarify the feature hierarchy within the network. Such studies might inform the design of custom architectures or loss functions optimized for subtle biometric distinctions. Our study uses ResNet-18 as the backbone of the recognition system. Future work could explore whether increasing the depth or complexity of the network architecture improves performance in separating MZ and NMZ iris. For example, architectures such as vision transformers (ViTs) [10] or deeper convolutional networks (e.g., ResNet-152) [13] could be tested.
@@ VIII CONCLUSION
In this study, we addressed the challenge of identifying MZ iris pairs from NMZ pairs — a problem that traditional iris recognition systems struggle with, given the subtle genetic similarities among MZ twins. By leveraging a Siamese network architecture and contrastive learning, we learned compact, discriminative representations capable of capturing nuanced patterns that define genetic similarity in the iris.
Our examination of different input conditions — original images, iris-only images, and non-iris-only images — revealed that the iris region itself is the critical component for determining genetic similarity. This understanding not only guided our model design but also offered insights into the underlying biometric signals at play.

### v3

@@ Abstract
This study presents the first automated classifier designed to determine whether a pair of iris images originates from monozygotic individuals, addressing a previously untackled problem in biometric recognition. In Daugman-style iris recognition, the textures of the left and right irises of the same person are traditionally considered as being as different as the irises of two unrelated persons. However, previous research indicates that humans can detect that two iris images are from different eyes of the same person, or eyes of monozygotic twins, with an accuracy of about 80%. In this work, we employ a Siamese network architecture and contrastive learning to categorize a pair of iris images as coming from monozygotic or non-monozygotic irises. This could potentially be applied, for example, as a fast, noninvasive test to determine if twins are monozygotic or non-monozygotic. We construct a dataset comprising both synthetic monozygotic pairs (images of different irises of the same individual) and natural monozygotic pairs (images of different images from persons who are identical twins), in addition to non-monozygotic pairs from unrelated individuals, ensuring a comprehensive evaluation of the model’s capabilities. To gain deeper insights into the learned representations, we train and analyze three variants of the model using (1) the original input images, (2) iris-only images (masking everything but the iris region), and (3) non-iris-only images (masking the iris region). This comparison reveals that both iris texture and surrounding ocular structure contain information useful for the model to classify the image pairs as monozygotic or non-monozygotic. Our approach achieves accuracy levels using the full iris image that exceed those previously reported for human classification of monozygotic iris pairs. [^1] Code is available on https://github.com/yongleyuan/mz-iris-pairing
@@ I INTRODUCTION
Iris recognition has emerged as a leading biometric identification technology due to the unique and stable patterns present in the iris [6]. These intricate patterns are formed randomly during fetal development [27]. Iris recognition technologies are able to identify the distinctiveness between irises from different individuals. High accuracy, combined with speed and non-intrusive nature, makes iris recognition ideal for applications in security, border control, and identity verification. For example, India has used iris as one of the biometrics in the unique identification system called “Aadhaar” [24].
While the principle that “no two irises are the same” supports the effectiveness of iris recognition technology [6], a notable limitation arises when dealing with monozygotic twins. Humans can often perceive subtle similarities in the iris patterns of monozygotic (“identical”) twins, yet iris biometric systems generally fail to detect these nuances [14]. Monozygotic twins share identical genetic material due to originating from a single fertilized egg, leading to similarities in many visual characteristics, including iris patterns [17]. To our knowledge, no automated classifier has been previously developed to specifically distinguish between monozygotic and non-monozygotic iris pairs.
The ability to accurately determine if two individuals are monozygotic twins using iris recognition is crucial for several reasons. First, it offers a cost-effective and non-invasive alternative for determining if twins are identical or fraternal. Second, investigating how iris recognition systems detect — or fail to detect — subtle texture similarities contributes to explainable artificial intelligence (AI), improving our understanding of the decision-making processes of these systems and increasing their trustworthiness.
The objective of this research is to develop a machine learning model capable of distinguishing between monozygotic and non-monozygotic twins. The contributions of this work include:
@@ III-D Dataset Construction
We know of no large datasets of MZ twin iris images available to the research community. This is a challenge for developing a training set for creating a model. However, the left and right irises of the same person can serve as a proxy for MZ irises. Therefore, we create a large training set of MZ iris image pairs by using left and right irises of the same person as a substitute for MZ iris image pairs coming from different persons of a MZ set of twins. The similarities and differences of these two types is a topic that merits further study when such data becomes more available. Nevertheless, results in this paper show that this approach to assembling a training set is effective.
@@
The training and validation data used in our work is available to the research community on the website of the Computer Vision Research Lab at Notre Dame [^2] https://cvrl.nd.edu/projects/data/. We do not have the rights to re-distribute the MZ twins iris images [18], but the list of image pairs used will be made available.
@@ IV METHODOLOGY
To implement the contrastive learning strategy, we utilize a Siamese network architecture [3]. While this architecture is established for similarity learning tasks, it is applied to this specific iris biometric challenge for the first time to our knowledge. A Siamese network consists of two identical sub-networks with shared weights and parameters, ensuring both inputs are processed in the same manner. Each sub-network functions as an encoder to map an input image to a fixed-dimensional embedding in the vector space. The similarity between the two embeddings is then evaluated using a distance metric, such as Euclidean distance. The Siamese network is particularly well-suited for similarity-based tasks, as it enforces consistency in the representation of both inputs and enables the network to focus on learning meaningful differences between MZ and NMZ pairs.
@@ V-B Model Backbone
We employ ResNet-18 as the backbone for our image encoder due to its proven effectiveness in various computer vision tasks [13]. ResNet-18 is a convolutional neural network (CNN) known for its residual connection, which addresses the vanishing gradient problem by allowing gradients to flow through shortcut connections that skip one or more layers. ResNet has multiple common variants, of which ResNet-18 is the shallowest. It is effective and easy to train. We explored deeper ResNet backbones (34, 50, 101) but found that they did not offer any clear increase in accuracy for this problem; see Appendix -B1.
@@ V-C2 Loss Function
where $\mathcal{L}$ is the loss, $y_{\text{pred}}$ is a binary label indicating whether the pair is similar/positive/MZ ($y_{\text{pred}}=1$) or dissimilar/negative/NMZ ($y_{\text{pred}}=0$), $D$ is the Euclidean distance in (2), and $m$ is the margin that defines the maximum acceptable distance between negative pairs. We set $m=1$ to normalize the distance between 0 and 1 after training so that we can easily recognize outliers if there are any. Since the distance should be normalized, we intuitively set the distance decision boundary/threshold to the midpoint: 0.5. We will discuss how this threshold affects the training and testing in Appendix -B2.
@@ V-D Model Training
The model is trained separately using three different types of inputs: (1) original iris images, (2) iris-only images, and (3) non-iris-only images. Comparing iris-only and non-iris-only results allows us to analyze how the parts of the image contribute to the task.
30% of the training data are randomly selected as validation data. These image pairs are not inputted in any forward passes. Instead, they are used after each forward pass to compute the evaluation metrics, which will be compared with those collected after inference.
We initialized the ResNet-18 backbone with the ImageNet pre-trained weights, which should already have some prior knowledge in distinguishing between different objects. As is standard for Siamese networks, we then fine-tuned the model with weights being shared across two identical backbones. Each time the model is trained with a batch size of 32 pairs. The loss function and the optimizer, as mentioned in Sec. V-C2 and V-C3, are kept the same throughout. All experiments are run 5 times, with each individual run having the same train-validation split and data batching.
@@ V-E Training Results
Across all three types of input images, the model demonstrated rapid convergence during training. Fig. 5 illustrates accuracy and loss during training. The training accuracy for all three input types quickly approaches 1, indicating that the model can distinguish between MZ and NMZ pairs in nearly all training samples after epoch 7. The minimal separation of lines suggests that the input type has a negligible impact on final training accuracy. The training loss follows a similar pattern of convergence, reaching near-zero around epoch 7. The iris-only image pairs show a faster decline in loss, suggesting that the iris information makes the model converge faster and easier under the same setting, compared to analyzing complete iris images and images without iris texture.
@@ V-F Testing Results
Table I presents the test outcomes for all three input types over five experimental runs. The model using original iris images achieves the best overall performance, exceeding the other two in terms of accuracy, recall, and F1 score. Although the non-iris-only input obtains the highest precision score, the advantage is marginal. Fig. 6 shows examples of false positive and false negative predictions for all three models.
@@ VI-A Results Analysis
The testing results indicate that the original input model, which utilizes the complete eye region, outperforms both the iris-only (masked) and the non-iris-only (inversely masked) models. At first, we anticipated the masked input model to excel by focusing exclusively on the iris’s unique biometric features. However, these results suggest that non-iris ocular features or contextual information present in the original images contribute positively to distinguishing MZ from NMZ pairs, at least under the given training conditions.
@@
  • Contextual Cues: The additional features surrounding the iris—such as sclera patterns, eyelid contours, or subtle lighting variations—help the model form more robust embeddings. Even though these cues are not traditionally considered primary biometric traits, they offer supplemental distinguishing information.
@@
As for the model with non-iris-only input, while it does not outperform the model with original input, its reasonable performance suggests that non-iris regions do aid in classification. It is surprising that non-iris-only input outperforms iris-only input. The numbers clearly indicate that the non-iris region provides valuable information for discrimination. It challenges our initial expectations about the primary location of critical signals.
@@ VI-B Pupil Dilation Effects
VI-C Comparison with prior work on human performance
@@ VII FUTURE WORKS
Systematic ablation studies examining which layers and filters contribute most to MZ/NMZ discrimination would clarify the feature hierarchy within the network. Such studies might inform the design of custom architectures or loss functions optimized for subtle biometric distinctions. Our study uses ResNet-18 as the backbone of the recognition system. Future work could explore whether increasing the depth or complexity of the network architecture improves performance in separating MZ and NMZ iris. For example, architectures such as vision transformers (ViTs) [10].
@@ VIII CONCLUSION
In this study, we present the first automated approach to address the challenge of identifying MZ iris pairs from NMZ pairs — a problem that traditional iris recognition systems struggle with, given the subtle genetic similarities among MZ twins. By leveraging a Siamese network architecture and contrastive learning, we learned compact, discriminative representations capable of capturing nuanced patterns that define genetic similarity in the iris.
Our examination of different input conditions — original images, iris-only images, and non-iris-only images — revealed that using the complete iris image yields the best performance. Interestingly, the non-iris region alone demonstrated better accuracy than the iris-only region. This finding challenges the notion of the iris region being the sole critical component for distinguishing MZ and NMZ iris pairs. The understanding of the interplay between different iris image regions offered new insights into the underlying biometric signals.